Find the value of m if 4 + 4m = 13 + m
Answer by marilynh(7)

4 + 4m = 13 + m

Subtract 4 from both sides of the equation:
4m = 9 + m

Subtract m from both sides:
3m = 9

Divide both sides by 3:
m = 3